# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v2

@@ Abstract
journal: Anonymized journal
orcid]https://orcid.org/$0$-0001-9894-2116 orcid]https://orcid.org/$0$-0002-9536-3087 orcid]https://orcid.org/$0$-0003-2730-7697 orcid]https://orcid.org/$0$-0001-5181-7115 orcid]https://orcid.org/$0$-0002-1421-3686 orcid]https://orcid.org/$0$-0003-1561-70736
@@ 1 Introduction
[…]
@@ 2.2 Usability studies on code completion tools
[…]
Another aspect investigated by HCI researchers focused on the users’ behavior while adopting CCTs. From this study, two main modalities emerged (Prather et al., 2024): in “acceleration” mode, the programmer is aware of their next steps and utilizes the tool to speed up the coding process. In contrast, in the “exploration” mode, the programmer is uncertain about how to proceed and employs the tool to investigate his options. Acceleration mode is usually preferred, as developers often use the tools to complete repetitive code that cannot be copy-pasted (Liang et al., 2024), and for recalling syntax they don’t remember (Xu et al., 2022). As a result, the more developers accept CCTs suggestions, the more they perceive themselves as productive (Bird et al., 2022).
Liang et al. (2024) conducted a study to understand the usability of CCTs (such as GitHub Copilot and Tabnine) by utilizing a structured survey distributed to 410 GitHub users who had engaged with CCTs. Questions covered participants’ motivations for using or not using AI assistants, the usability issues they encountered, and the strategies they used to optimize the tool’s output. The survey revealed that developers are most motivated to use AI programming assistants because they help developers reduce keystrokes, finish programming tasks quickly, and recall syntax. They also found that developers often do not use CCTs because these tools do not generate code that addresses certain functional or non-functional requirements, and because developers have trouble controlling the tool to generate the desired output.
Another important requirement that emerged from user studies is the need for CCTs to learn from feedback and adapt their behavior to developers’ styles and project needs. For instance, Zhang et al. (2023) highlight that programmers would like to customize the shortcuts and to set suggestions length and frequency. The literature also highlights the need for explanations (such as reporting the source or linking to documentation) for additional context on the generated code (Liang et al., 2024; Xu et al., 2022).
The proactivity of CCTs has also been investigated, with research encouraging its implementation (Sergeyuk et al., 2024). Vaithilingam et al. (2023) suggested that automatic inline gray-text suggestions, as opposed to traditional lightbulb interfaces, significantly improve discoverability and usability of CCTs. The authors established five design principles to guide future AI-driven code suggestion interfaces:
@@ 2.3 The importance of mental models
A mental model is generally a mental picture users have of how a certain system works (Norman, 1983, 2013). It includes a user’s assumptions, beliefs, and expectations while interacting with a system that will lead to their actions, interactions, and problem-solving behavior. Mental models also change over time, might contain errors, and are constrained by the user’s prior experiences or technical background. Multiple mental models of a single object may be held by one person to reflect its many functions, and two persons will not always have the same mental model. If mental models are accurate and considered in the design of a system, users can anticipate system responses, feel confident during the interaction, and adapt quickly to new features. On the other hand, gaps between the users’ mental model of the system’s functionality and its actual behavior can lead users to confusion, frustration, and errors, which will reduce their productivity and satisfaction (Doi, 2021).
@@
The need to understand mental models in the domain of AI-based CCTs is multifaceted, and different aspects can benefit from a deep understanding of users’ mental models. The first is interaction with CCTs since users might predict the tool’s behavior and know when a suggestion will be useful and when it could mislead. For instance, a tool designed considering users’ mental models might allow developers to know when the AI might make contextually relevant suggestions and when it can blunder, and so on, to keep the coding flow in hand and avoid wasting time for distraction. Another aspect that can benefit regards trust. CCTs can be largely obscure to those operating outside the normal programming logic. It has been proven that transparency and interpretability are essential for establishing user trust, and understanding the user’s mental model is essential to making users comfortable when adopting the tools (Desolda et al., 2024). This helps create trust in a tool by aligning design features with users’ mental models, making developers more likely to adopt it into their daily workflow. Another aspect that can benefit from the mental model is cognitive load: CCTs are meant to reduce repetitive tasks and cognitive load; thus, users can concentrate on more difficult problem-solving jobs. However, if users cannot intuitively understand how or why these suggestions are being made, the cognitive load can actually rise rather than fall. Refining these tools to become less disjointed in relation to user mental models means we can improve the flow of work, minimize interruptions, and ultimately help users be more productive. The last aspect is the possible support to learning: especially for novice programmers, developers can recognize common coding patterns and problem-solving techniques, enhancing both tool competency and coding proficiency. The overall improvements in systems that stem from following users’ mental models throughout the design are also a potent driver for increasing systems’ adoption and acceptability (Sergeyuk et al., 2025). A recent study by Wang et al. (2023) highlights that 54% of the developers recruited for their study are eager for better CCTs.
@@
Exploring the users’ mental models is a complex process since these models are typically abstract, partially formed, or even subconscious. Without elicitation studies, bridging this gap is an arduous task, and they are the ones that prescribe structured methodologies to help yield, analyze, and interpret users’ cognitive models regarding system functionality. The techniques for eliciting mental models include design workshops (Jacko, 2012), interviews (Rogers et al., 2023), concept mapping (Jacko, 2012), think-aloud protocols (Lewis, 1982), and scenario-based design (Rosson and Carroll, 2012). These techniques allow researchers to design intricate representations of users’ mental models of system behavior, what they expect the system to do a priori, and how they can expect to interact with, for example, CCTs. Such insights inform tool design and can help bolster our understanding of HCI within AI-augmented coding domains towards supporting frameworks and guidelines for improved user experiences in such environments and beyond.
@@ 3.1 Code completion tools analysis
The analysis of the remaining 13 CCTs providing a User Interface (UI) was guided by two main aspects: the code suggestion itself and the explanation, if any, of why that code was suggested. After familiarizing ourselves with the CCTs, we analyzed users’ interaction with them across different dimensions. To simplify the analysis and the reporting of the results, we used the 5W model by Lasswell (1974), which is adopted in various fields, such as journalism and customer analysis, and more generally in problem-solving, to analyze the complete story of a fact. A similar approach is employed in the field of HCI, where the 5W model is utilized for mental model elicitation, and comprehensive user analysis (Jacko, 2012; Desolda et al., 2017; Duric et al., 2002). Each dimension of the interaction with the tool was addressed in terms of the question of the 5W model. In this phase, the 5W model guides the analysis of the state of the art of CCTs. The questions used in this phase are similar to those presented in Section˜4.3, which are used to elicit users’ mental model: this choice allows the comparison of existing CCTs with users’ expectations. This choice allows to leverage existing tools (identified during this analysis) as design inspirations during prototype development (see Section˜7), since it allows the potential identification of tools that satisfy users’ expectations (elicited in the co-design workshops, see Section˜4). The analysis of the tools with respect to the code suggestion was done by answering these questions:
@@
Intellicode (Svyatkovskiy et al., 2020) lines proactive grey text curent file inline - - -
Codex (Chen et al., 2021) lines manual text chat history right side answer prompt chatbot
Copilot (Friedman, 2021) lines proactive grey text current file inline answer prompt chatbot
TravTrans (Kim et al., 2021) token proactive dropdown current file inline - - -
StarCoder (Li et al., 2023) lines proactive grey text current file inline - - -
Codegeex (Zheng et al., 2023) lines proactive grey text current file inline answer, description prompt, click chatbot
CodeWhisperer (Amazon, 2024) lines hybrid grey text current file inline source click sidebar
Codeium (Codeium, 2024) lines proactive grey text current file inline improvements, reasons click chatbot
Replit (Replit Inc., 2024) lines proactive grey text current file inline description click pointer
Cody (Sourcegraph, 2025) lines proactive grey text repository inline answer click, prompt chatbot
Tabnine (Tabnine, 2025) lines proactive grey text current file inline answer prompt chatbot
Xcode (Apple Inc., 2024) lines proactive grey text current file inline - - -
Blackbox (Blackbox AI, 2024) lines proactive grey text current file inline improvement click sidebar
@@ 3.2 Code suggestions
An important characteristic affecting the interaction with CCTs is the timing to provide suggestions (when). The analysis of the existing tools revealed two main approaches: proactive and manual. Proactive CCTs provide suggestions at a certain time without requiring a user action to trigger the completion. This type of design can be either implemented in an intrusive manner (by suggesting a completion at any time) or discreetly (by generating completions only when the context is sufficient). The manual approach, in contrast, requires the user to perform a specific action to visualize the completion. However, manual activation of CCTs is not very popular, as it is adopted in only two tools. A hybrid timing approach is also implemented, even if only in the case of CodeWhisperer, as completions are proactive but can be forced through a shortcut (Amazon, 2024).
Another aspect is the content of CCTs suggestions (what is provided as a suggestion?). Two main modalities emerged from our analysis: single token suggestions (less popular, as it was proposed only with the TravTrans prototype (Kim et al., 2021)) and multiple line completions, which are widely adopted. While the first solution is quite simple, the second one results in a feature that is hard to control and assess since the length of the suggestions may vary depending on the context and LLM characteristics. As a matter of fact, multi-line suggestion length can range from short code blocks (e.g., for loops, switches, if-else) to entire functions or classes.
The suggestion’s location (where) in the UI is also worth exploring. A predominant approach is to put completions directly inside the code (inline suggestions). The other solution instead consists of providing the suggestion in a separate space with respect to the coding area. Codex (Chen et al., 2021) is the only tool that adopts the latter modality.
@@
Finally, it is worth investigating the customizability of CCTs, as the relevance and helpfulness of suggestions strongly depend on the type of user receiving them (how to customize). Context-aware suggestions may consider the user expertise, the specific formatting/adherence to coding standards, and the functional and non-functional requirements of specific software development scenarios to provide the most accurate code snippets possible. The strategies adopted for this aim result in three categories: personalization based on the context of the current file, personalization based on a code repository, and personalization based on chat history. The first category is the most popular, while Cody (Sourcegraph, 2025) is the only tool that managed to include a code repository for more relevant completions. Nevertheless, this last approach seems to fail for big repositories. Codex provides more aware suggestions by processing the conversation between the programmer and the chatbot as context.
@@ 3.3 Explanations of the suggested code
The location in the UI (where) of code explanations is an important aspect of CCTs, as it influences other interaction characteristics. Three main approaches are observed: explanations are either delivered inside a chatbot, or shown in a sidebar, or finally displayed inside the coding space. Tabnine (Tabnine, 2025), Cody (Sourcegraph, 2025), CodeGeeX (Zheng et al., 2023), Copilot (Friedman, 2021), Codeium (Codeium, 2024), and Codex (Chen et al., 2021) implement the first approach, providing chat-based explanations. In contrast, CodeWhisperer (Amazon, 2024) and Blackbox (Blackbox AI, 2024) report code explanations in a sidebar. In this case, the user is not allowed of submitting prompts for additional details, but such an approach delivers code explanations quicker, as the user does not have to think about a prompt in order to receive explanations. Finally, Replit (Replit Inc., 2024) follows a unique approach, showing explanations inside a tooltip at the mouse pointer’s location (which means the explanations appear within the coding space).
@@ 4 Methodology
This section details the methodology that was employed in this study. Leveraging the 5W model by Lasswell (1974), we conducted co-design workshops that aimed at eliciting users’ mental models and expectations for CCTs.
@@ 4. RQ4:
To answers these questions, we followed an elicitation study methodology conducted through design workshops. We have chosen this technique because, particularly when performed in workshop formats, it facilitates in-depth insights into users’ thought processes, preferences, expectations, needs, and perceptions about complex systems (similar to what has been done by Desolda et al. (2017); Marquardt et al. (2012); Voida et al. (2005)). Moreover, workshops stimulate participants to collaboratively propose, discuss, and model their ideas on code completion tools, offering rich qualitative data suited to the study’s objectives. To structure the workshop discussions, we adopted a scenario-based design approach. This solution helps participants provide their answers while referring to a realistic situation (Rosson and Carroll, 2012).

### v3

@@ Abstract
journal: Int. J. Hum.-Comp. Studies
@@
{highlights}
Developers expressed diverse preferences on when and how code suggestions should be triggered; this highlights the need for both user control and context-sensitive automation.
The position to visualize code suggestion and explanation can be inline for short completions; sidebar, pop-up, or chatbot for longer or more complex suggestions.
Most users want to receive minimal code suggestions and then request details on demand, including support for varying from code statements to whole functions or files.
Explanations should be concise, contextual, and user-triggered (via click or shortcut), with links to documentation or example code.
Developers prefer to customize activation timing, visualization style, explanation detail, and code style according to their personal preferences and project needs.
@@ 1 Introduction
While the technical aspects of AI’s role in code generation are well-researched (for example, algorithms and AI models used by such tools (Izadi et al., 2024; Husein et al., 2024)), HCAI systems for code completions are still underresearched, exposing developers (and their software) to various risks. Some recent studies confirm these symptoms: only 40% of CCT suggestions are accepted the first time they are proposed to developers, mainly because developers cannot control context or granularity (Liang et al., 2024). Other evidence about the misalignment between developers’ mental models and CCTs regards the developers’ need for finer-grained suggestions and trustworthy explanations (Sergeyuk et al., 2025; Wang et al., 2023), typically not provided in CCTs, as well as the different preferences about inline and sidebar layouts (Vaithilingam et al., 2023; Mărășoiu et al., 2015). Although some studies explored coders’ expectations when dealing with CCTs (Sergeyuk et al., 2025; Wang et al., 2023), the understanding of the cognitive and interaction dynamics that developers face upon adopting CCTs is limited. To the best of our knowledge, no existing studies empirically map the mental models of developers on the interaction with CCTs, models that might determine when code suggestions and explanations are trusted, understood, or disrupted. This aspect limits both theoretical understanding and practical guidance for interaction designers of CCTs. Understanding developers’ mental models can thus lead to developing an IDEs that aligns with user requirements, fosters user satisfaction, and improves coding productivity while promoting their adoption (Sergeyuk et al., 2025).
@@ 2.2 Usability studies on code completion tools
The usability of CCTs has been a research focus for quite some time, aiming to make these tools more intuitive and helpful for developers. One of the first attempts was carried out by Mără˙ (2015), which empirically investigated how professional developers interact with CCTs. They analyzed the behaviors, intentions, and obstacles programmers encounter when employing these instruments. The study results revealed that code completion is primarily employed to speed up the coding process and guarantee accuracy. Besides, developers frequently use it as real-time feedback to identify mistakes.
Another aspect investigated by HCI researchers focused on the users’ behavior while adopting CCTs. From this study, two main modalities emerged (Prather2024Its): in “acceleration” mode, the programmer is aware of their next steps and utilizes the tool to speed up the coding process. In contrast, in the “exploration” mode, the programmer is uncertain about how to proceed and employs the tool to investigate his options. Acceleration mode is usually preferred, as developers often use the tools to complete repetitive code that cannot be copy-pasted (Liang2024LargeScale), and for recalling syntax they don’t remember (Xu2022IDE). As a result, the more developers accept CCTs suggestions, the more they perceive themselves as productive (Bird2022Taking).
Liang2024LargeScale conducted a study to understand the usability of CCTs (such as GitHub Copilot and Tabnine) by utilizing a structured survey distributed to 410 GitHub users who had engaged with CCTs. Questions covered participants’ motivations for using or not using AI assistants, the usability issues they encountered, and the strategies they used to optimize the tool’s output. The survey revealed that developers are most motivated to use AI programming assistants because they help developers reduce keystrokes, finish programming tasks quickly, and recall syntax. They also found that developers often do not use CCTs because these tools do not generate code that addresses certain functional or non-functional requirements, and because developers have trouble controlling the tool to generate the desired output.
Another important requirement that emerged from user studies is the need for CCTs to learn from feedback and adapt their behavior to developers’ styles and project needs. For instance, Zhang2023Demystifying highlight that programmers would like to customize the shortcuts and to set suggestions length and frequency. The literature also highlights the need for explanations (such as reporting the source or linking to documentation) for additional context on the generated code (Liang2024LargeScale; Xu2022IDE).
The proactivity of CCTs has also been investigated, with research encouraging its implementation (Sergeyuk2024IDE). Vaithilingam2023More suggested that automatic inline gray-text suggestions, as opposed to traditional lightbulb interfaces, significantly improve discoverability and usability of CCTs. The authors established five design principles to guide future AI-driven code suggestion interfaces:
@@ 2.3 The importance of mental models
A mental model is generally a mental picture users have of how a certain system works (Norman1983Observations; Norman2013Design). It includes a user’s assumptions, beliefs, and expectations while interacting with a system that will lead to their actions, interactions, and problem-solving behavior. Mental models also change over time, might contain errors, and are constrained by the user’s prior experiences or technical background. Multiple mental models of a single object may be held by one person to reflect its many functions, and two persons will not always have the same mental model. If mental models are accurate and considered in the design of a system, users can anticipate system responses, feel confident during the interaction, and adapt quickly to new features. On the other hand, gaps between the users’ mental model of the system’s functionality and its actual behavior can lead users to confusion, frustration, and errors, which will reduce their productivity and satisfaction (190/sym130.33050795).
@@
The need to understand mental models in the domain of AI-based CCTs is multifaceted, and different aspects can benefit from a deep understanding of users’ mental models. The first is interaction with CCTs since users might predict the tool’s behavior and know when a suggestion will be useful and when it could mislead. For instance, a tool designed considering users’ mental models might allow developers to know when the AI might make contextually relevant suggestions and when it can blunder, and so on, to keep the coding flow in hand and avoid wasting time for distraction. Another aspect that can benefit regards trust. CCTs can be largely obscure to those operating outside the normal programming logic. It has been proven that transparency and interpretability are essential for establishing user trust, and understanding the user’s mental model is essential to making users comfortable when adopting the tools (RN1500). This helps create trust in a tool by aligning design features with users’ mental models, making developers more likely to adopt it into their daily workflow. Another aspect that can benefit from the mental model is cognitive load: CCTs are meant to reduce repetitive tasks and cognitive load; thus, users can concentrate on more difficult problem-solving jobs. However, if users cannot intuitively understand how or why these suggestions are being made, the cognitive load can actually rise rather than fall. Refining these tools to become less disjointed in relation to user mental models means we can improve the flow of work, minimize interruptions, and ultimately help users be more productive. The last aspect is the possible support to learning: especially for novice programmers, developers can recognize common coding patterns and problem-solving techniques, enhancing both tool competency and coding proficiency. The overall improvements in systems that stem from following users’ mental models throughout the design are also a potent driver for increasing systems’ adoption and acceptability (Sergeyuk et al., 2025). A recent study by Wang et al. (2023) highlights that 54% of the developers recruited for their study are eager for better CCTs.
@@
Exploring the users’ mental models is a complex process since these models are typically abstract, partially formed, or even subconscious. Without elicitation studies, bridging this gap is an arduous task, and they are the ones that prescribe structured methodologies to help yield, analyze, and interpret users’ cognitive models regarding system functionality. The techniques for eliciting mental models include design workshops (Jacko2012Human), interviews (Rogers2023Interaction), concept mapping (Jacko2012Human), think-aloud protocols (Lewis1982Using), and scenario-based design (Rosson2012ScenarioBased). These techniques allow researchers to design intricate representations of users’ mental models of system behavior, what they expect the system to do a priori, and how they can expect to interact with, for example, CCTs. Such insights inform tool design and can help bolster our understanding of HCI within AI-augmented coding domains towards supporting frameworks and guidelines for improved user experiences in such environments and beyond.
@@ 3.1 Code completion tools analysis
The analysis of the remaining 13 CCTs providing a User Interface (UI) was guided by two main aspects: the code suggestion itself and the explanation, if any, of why that code was suggested. After familiarizing ourselves with the CCTs, we analyzed users’ interaction with them across different dimensions. To simplify the analysis and the reporting of the results, we used the 5W model by Lasswell1974Structure, which is adopted in various fields, such as journalism and customer analysis, and more generally in problem-solving, to analyze the complete story of a fact. A similar approach is employed in the field of HCI, where the 5W model is utilized for mental model elicitation, and comprehensive user analysis (Jacko2012Human; Desolda2017Empowering; Duric2002Integrating). Each dimension of the interaction with the tool was addressed in terms of the question of the 5W model. In this phase, the 5W model guides the analysis of the state of the art of CCTs. The questions used in this phase are similar to those presented in Section˜4.3, which are used to elicit users’ mental model: this choice allows the comparison of existing CCTs with users’ expectations. This choice allows to leverage existing tools (identified during this analysis) as design inspirations during prototype development (see Section˜7), since it allows the potential identification of tools that satisfy users’ expectations (elicited in the co-design workshops, see Section˜4). The analysis of the tools with respect to the code suggestion was done by answering these questions:
@@
Intellicode (Svyatkovskiy2020IntelliCode) lines proactive grey text curent file inline - - -
Codex (Chen2021Evaluating) lines manual text chat history right side answer prompt chatbot
Copilot (Friedman2021Introducing) lines proactive grey text current file inline answer prompt chatbot
TravTrans (Kim2021Code) token proactive dropdown current file inline - - -
StarCoder (li2023starcoder) lines proactive grey text current file inline - - -
Codegeex (Zheng2023CodeGeeX) lines proactive grey text current file inline answer, description prompt, click chatbot
CodeWhisperer (What) lines hybrid grey text current file inline source click sidebar
Codeium (Codeium) lines proactive grey text current file inline improvements, reasons click chatbot
Replit (Replit) lines proactive grey text current file inline description click pointer
Cody (Cody) lines proactive grey text repository inline answer click, prompt chatbot
Tabnine (Tabnine) lines proactive grey text current file inline answer prompt chatbot
Xcode (Xcode) lines proactive grey text current file inline - - -
Blackbox (Chat) lines proactive grey text current file inline improvement click sidebar
@@ 3.2 Code suggestions
An important characteristic affecting the interaction with CCTs is the timing to provide suggestions (when). The analysis of the existing tools revealed two main approaches: proactive and manual. Proactive CCTs provide suggestions at a certain time without requiring a user action to trigger the completion. This type of design can be either implemented in an intrusive manner (by suggesting a completion at any time) or discreetly (by generating completions only when the context is sufficient). The manual approach, in contrast, requires the user to perform a specific action to visualize the completion. However, manual activation of CCTs is not very popular, as it is adopted in only two tools. A hybrid timing approach is also implemented, even if only in the case of CodeWhisperer, as completions are proactive but can be forced through a shortcut (What).
Another aspect is the content of CCTs suggestions (what is provided as a suggestion?). Two main modalities emerged from our analysis: single token suggestions (less popular, as it was proposed only with the TravTrans prototype (Kim2021Code)) and multiple line completions, which are widely adopted. While the first solution is quite simple, the second one results in a feature that is hard to control and assess since the length of the suggestions may vary depending on the context and LLM characteristics. As a matter of fact, multi-line suggestion length can range from short code blocks (e.g., for loops, switches, if-else) to entire functions or classes.
The suggestion’s location (where) in the UI is also worth exploring. A predominant approach is to put completions directly inside the code (inline suggestions). The other solution instead consists of providing the suggestion in a separate space with respect to the coding area. Codex (Chen2021Evaluating) is the only tool that adopts the latter modality.
@@
Finally, it is worth investigating the customizability of CCTs, as the relevance and helpfulness of suggestions strongly depend on the type of user receiving them (how to customize). Context-aware suggestions may consider the user expertise, the specific formatting/adherence to coding standards, and the functional and non-functional requirements of specific software development scenarios to provide the most accurate code snippets possible. The strategies adopted for this aim result in three categories: personalization based on the context of the current file, personalization based on a code repository, and personalization based on chat history. The first category is the most popular, while Cody (Cody) is the only tool that managed to include a code repository for more relevant completions. Nevertheless, this last approach seems to fail for big repositories. Codex provides more aware suggestions by processing the conversation between the programmer and the chatbot as context.
@@ 3.3 Explanations of the suggested code
The location in the UI (where) of code explanations is an important aspect of CCTs, as it influences other interaction characteristics. Three main approaches are observed: explanations are either delivered inside a chatbot, or shown in a sidebar, or finally displayed inside the coding space. Tabnine (Tabnine), Cody (Cody), CodeGeeX (Zheng2023CodeGeeX), Copilot (Friedman2021Introducing), Codeium (Codeium), and Codex (Chen2021Evaluating) implement the first approach, providing chat-based explanations. In contrast, CodeWhisperer (What) and Blackbox (Chat) report code explanations in a sidebar. In this case, the user is not allowed of submitting prompts for additional details, but such an approach delivers code explanations quicker, as the user does not have to think about a prompt in order to receive explanations. Finally, Replit (Replit) follows a unique approach, showing explanations inside a tooltip at the mouse pointer’s location (which means the explanations appear within the coding space).
@@ 4 Methodology
This section details the methodology that was employed in this study. Leveraging the 5W model by Lasswell1974Structure, we conducted co-design workshops that aimed at eliciting users’ mental models and expectations for CCTs.
@@ 4. RQ4:
To answers these questions, we followed an elicitation study methodology conducted through design workshops. We have chosen this technique because, particularly when performed in workshop formats, it facilitates in-depth insights into users’ thought processes, preferences, expectations, needs, and perceptions about complex systems (similar to what has been done by Desolda2017Empowering; Marquardt2012Crossdevice; Voida2005Study). Moreover, workshops stimulate participants to collaboratively propose, discuss, and model their ideas on code completion tools, offering rich qualitative data suited to the study’s objectives. To structure the workshop discussions, we adopted a scenario-based design approach. This solution helps participants provide their answers while referring to a realistic situation (Rosson2012ScenarioBased).